40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4007     Entry time: Fri Dec 3 05:21:11 2010
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: locked the laser to the cavity 

I succeeded in locking the end green laser to X arm with the new ETM.

Though the lock is still not so stable compared to the previous locking with the old ETM. Also the beam centering is quite bad now.

So I will keep working on the end green lock a little bit more.

Once the lock gets improved and becomes reasonably stiff, we will move onto the corner PLL experiment.

 

(to do)

 - beam centering on ITMX

 - check the mode matching

 - revise the control servo

ELOG V3.1.3-